If five times the lesser of two consecutive integers is added to three times the greater, the result is 59. Find the integers.
Answer by checkley77(12844)

Let x & x+1 be the 2 integers.
5x+3(x+1)=59
5x+3x+3=59
8x=59-3
8x=56
x=56/8
x=7 ans for the first integer.
7+1=8 ans. for the second integer.
Proof:
5*7+3*8=59
35+24=59
59=59